Discuss the thermolysis methods for generation of the free radicals.
1
Expert's answer
2019-03-25T07:31:05-0400

The most obviuos method for radical generation is to heat molecules until the bond begins to break.

If a molecule or ion has one ralatively weak bond, precise homolysis may occur on heating:


"A-B \xrightarrow{heat} A\cdot + B\cdot"

An example is the thermolysis of (i) poroxides or of (ii) azo-compounds:


"(i) R-O-O-R \xrightarrow{heat} 2R-O \cdot"

"(ii) R-N=N-R \xrightarrow{heat} 2R \cdot + N_2"